OpenStudy (anonymous):

Factor using the quadratic formula: x^2+x-30

OpenStudy (solomonzelman):

\[30=(-5) \times 6\]

OpenStudy (neer2890):

Roots of quadratic equation \[ax ^{2}+bx+c=0\] \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

Ah, thats what I was doing wrong. For some reason I was plugging the variables into the equation, not the coefficients. That's why it didn't make any sense. Thanks! I think I can get it from here.

OpenStudy (neer2890):

you're welcome...:)

OpenStudy (anonymous):

Hmm I started working through, and I came up with this: \[\frac{ -1\pm \sqrt{120} }{ 2 }\] Now I'm not quite sure what to do with it, am I at least on the right track?

OpenStudy (neer2890):

\[x=\frac{ -1\pm \sqrt{1-4*1*(-30)} }{ 2 }=\frac{ -1\pm \sqrt{1+120} }{ 2 }\]

OpenStudy (neer2890):

here, a=1 b=1 and c=-30 just plug in the values and find the value of x.

OpenStudy (anonymous):

Bah, missed a step. Thanks again!

OpenStudy (neer2890):

don't mention.

OpenStudy (anonymous):

(x-5)(x+6) :)